OpenStudy (priyar):

For natural numbers m,n if (1-y)^m (1+y)^n = 1+ay+by^2 +..... and a=b= 10, then (m,n) is ?

OpenStudy (priyar):

@Jaynator495 @jigglypuff314 @TheSmartOne @Luigi0210 @robtobey @agent0smith

OpenStudy (priyar):

i got 2 eqs nC1-mC1 = 10 and nC2 +mC2 = 10 ...are these correct?

imqwerty (imqwerty):

\(\left[~^mC_0(-y)^0 +~^mC_1(-y)^1+....~^mC_m(-y)^m \right] \left[~^nC_0(y)^0+~^nC_1y^1+...+~^nC_n(y)^n\right]\) you will get the term of format ay only when you multiply something like-> 1 \times y make such cases and then add al cases where ay thing is forming same for by^2

OpenStudy (priyar):

ya i did like this for the first two terms and got the abv.. 2 eqs are they correct?

imqwerty (imqwerty):

i am checking the 2nd one 1st one is correct

OpenStudy (priyar):

ok..

OpenStudy (priyar):

@imqwerty?

imqwerty (imqwerty):

sorry there was some network error..

OpenStudy (priyar):

oh..its ok

imqwerty (imqwerty):

okay the 2nd equation is not correct you can get a y^2 term on multiplying 2 y terms

OpenStudy (priyar):

oh ok...got it!

imqwerty (imqwerty):

:)

OpenStudy (priyar):

Thanks!

OpenStudy (priyar):

can you help in the other one also?